# Centralized Reward Agent for Knowledge Sharing and Transfer in Multi-Task Reinforcement Learning

### v2

@@ 4.1.3 Policy Agents with Knowledge Rewards
Each policy agent $\mathcal{A}^{pol}_{i}$ interacts with its specific environment and stores the experiences in its corresponding replay buffer $\mathcal{D}_{i}$. Policy agents receive two types of rewards: the environmental reward $r^{env}_{i}$ from their respective task $T_{i}$ and the knowledge reward $r^{knw}$ from CRA. The augmented reward is given by:
@@ 4.2 Information Synchronization of Policy Agents
In the CenRA framework, the information provided by different tasks may exhibit significant disparities, potentially leading to an imbalance in knowledge extraction and distribution. In this section, we introduce an information synchronization mechanism for CenRA to maintain a balanced manner from the perspective of the entire system. Specifically, this mechanism is implemented by controlling the quantity of samples that CRA retrieves from each task’s replay buffer $\mathcal{D}_{i}$ by a sampling weight $\bm{w}$. We mainly consider two aspects: the similarity among tasks and the real-time learning performance of the policy agents.
Similarity Weight is derived from the similarity among tasks, enabling the CRA to focus on relatively outlier tasks. To simplify computation, we use the hidden layers extracted from each policy agent’s neural network encoders to represent the tasks’ features. To reduce randomness, we average the hidden features of the most recent $K$ steps. We adopt a cross-attention mechanism, which is widely used in neural networks, to calculate the similarity weight Vaswani et al., (2017). Specifically, for task $T_{i}$, let $\bm{H}_{i}$ denote the averaged hidden feature vector, which serves as the key, and the centroid of all tasks $\bm{c}$ acts as the query. Then, the similarity $s_{i}$ of task $T_{i}$ to the centroid of the task cluster is calculated as:
@@
where $D$ is the dimension of the hidden feature to prevent gradient vanishing or exploding. A larger $s_{i}$ indicates a greater similarity between $T_{i}$ and the centroid. Given our assumption is that the tasks farther from the centroid require more attention, the similarity weight is defined as $\bm{w}^{sim}=\text{Softmax}\big{(}[1/s_{1},1/s_{2},\dots,1/s_{N}]\big{)}$.
@@
Lastly, the final sampling weight $\bm{w}$ combines the similarity weight and the performance weight as $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$, where $\alpha$ is a hyperparameter to balance the two aspects. The CRA samples from each replay buffer $\mathcal{D}_{i}$ according to $\bm{w}$, ensuring a balanced and effective knowledge extraction and learning.
@@ 4.3 Overall Framework
The learned CRA acts as a robust knowledge pool, which is able to support new tasks as an additional module by transferring knowledge through auxiliary reward signals. This is particularly beneficial in sparse-reward environments, as the knowledge rewards can guide the policy agents toward the correct direction and reduce the exploration burden. Additionally, the CRA can be further optimized alongside new tasks in a continuous learning scheme that enhances adaptability and effectiveness in dynamic settings.
@@
3:        $(s_{t},a_{t},r^{env}_{t},s_{t+1},a_{t+1})\sim\text{Interact}(\mathcal{A}^{pol}_{i},T_{i})$triangleright Interact and collect one transition
4:        $r^{knw}_{t}\sim\mathcal{A}^{rwd}(s_{t},a_{t})$triangleright Sample an off-policy knowledge reward
@@
6:        $\mathcal{D}_{i}\leftarrow\mathcal{D}_{i}\cup\{(s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})\}$triangleright Store the transition in corresponding $\mathcal{D}_{i}$
7:        Update policy agent $\mathcal{A}^{pol}_{i}$triangleright Update $\mathcal{A}^{pol}_{i}$ using backbone RL algorithm
@@
9:     $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$triangleright Calculate sampling weight
10:     $\{s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i}\}_{\mathcal{B}}\sim\mathcal{D}|\bm{w}$triangleright Draw samples based on the sampling weight
@@ 5 Experiments
We conduct experiments in four MTRL domains as shown in Figure 3: the widely used Meta-World benchmark (including ML10 with 10 tasks and ML50 with 50 tasks) (Yu et al., 2020b, ), 2DMaze, 3DPickup (Chevalier-Boisvert et al.,, 2024), and MujocoCar (Ji et al.,, 2023). All tasks, including those in Meta-World, are crafted to provide sparse environmental rewards, where the agent receives a reward of $1$ only upon successful completion of the final objective, and 00 otherwise. The detailed task configurations are provided in Appendix A.
@@
Figure 3: Environments with multiple tasks. (a) Meta-World: two sparse-reward versions are used: ML10-sparse and ML50-sparse, including diverse robotic manipulation tasks. (b) 2DMaze: 2D maze tasks where the agent must pick up a key and then pass through a door to exit. (b) 3DPickup: 3D maze tasks where the agent aims to navigate to and pick up different target objects at different locations. (c) MujocoCar: mujoco-based race car aims to navigate to different specified areas.
@@ 5.1 Comparative Evaluation in MTRL
We benchmark CenRA against several state-of-the-art baselines: (a) the backbone RL algorithms of the policy agents: DQN (Mnih et al.,, 2015) for discrete control tasks and SAC (Haarnoja et al., 2018b, ) for continuous control tasks; (b) the ReLara algorithm (Ma et al., 2024a, ), which can be regarded as a decentralized variant of CenRA, where each policy agent is paired with a separate reward agent, without cross-task information sharing; (c) the Policy Optimization and Policy Correction (PiCor) algorithm (Bai et al.,, 2023); (d) the Multi-Critic Actor Learning (MCAL) algorithm (Mysore et al.,, 2022); (e) the Parameter-compositional MTRL (PaCo) algorithm (Sun et al.,, 2022); and (f) the MTRL with Soft Modularization (SoftModule) (Yang et al.,, 2020). They are implemented by either the CleanRL library (Huang et al.,, 2022) or official codebases. Each task is trained with 10 different random seeds, and the average results are reported.
@@
5.2 What Has the Centralized Reward Agent Learned?
(a) Maze #1
(b) Maze #2
(c) Maze #3
(d) Maze #4
Figure 5: Visualization of the action yielding the maximum knowledge rewards in the four 2DMaze tasks.
@@ 5.3 Knowledge Transfer to New Tasks
In this section, we assess the CRA’s ability to transfer previously learned knowledge to unseen tasks. Specifically, we continue using the trained CRA model in Section 5.1, while initializing new policy agents to tackle new tasks from the same domain. These include 5 test tasks for ML10-sparse and ML50-sparse, and 1 test task for each of the remaining domains, none of which were encountered during the initial training. For the CenRA, we explore two scenarios: (1) the CRA continues to be optimized in collaboration with the new policy agent (CenRA w/ learning); and (2) only the policy agent is updated while the CRA remains fixed, relying only on its previously acquired knowledge (CenRA w/o learning). We compare the two settings against the backbone algorithms and ReLara. In ReLara, the reward agent is trained anew without pre-learned knowledge. The results are presented in Figure 6 and Table 2.
@@
Figure 6: Comparison of the learning performance of CenRA with the baselines in new tasks.
@@
(a) Before getting the key.
(b) After getting the key.
Figure 7: Actions yielding the highest knowledge rewards at each state in a new 2DMaze task.
To further demonstrate CenRA’s transferability, we select the 2DMaze environment to visualize the knowledge provided by CRA when facing an unseen task. As shown in Figure 7, we plot the directions of actions that yield the maximum knowledge reward at each position, categorized into two scenarios: before and after obtaining the key. While some guidance in peripheral regions may appear slightly misaligned, most states receive reasonable rewards that align with human understanding. This demonstrates the effectiveness of knowledge transfer, and with such dense rewards, the agent’s adaptation to new tasks is able to be well supported.
@@ 5.4 Effect of Sampling Weight
We conducted experiments to understand the effects of the information synchronization mechanism in the CenRA. Specifically, we compare the full CenRA model against three variants: (a) CenRA without the similarity weight $\bm{w}^{sim}$, (b) CenRA without the performance weight $\bm{w}^{per}$, and (c) CenRA without the entire sampling weight. To better illustrate the differences among tasks and highlight the role of sampling weights in task coordination and synchronization, we select the four-task domains, i.e., 2DMaze, 3DPickup, and MujocoCar. The comparison results are shown in Table 3.
Algo. 2DMaze Var. $\downarrow$ (×$\times 10^{-2}$)
Maze #1 Maze #2 Maze #3 Maze #4
CenRA 0.89 $\pm$ 0.03 0.91 $\pm$ 0.02 0.92 $\pm$ 0.02 0.93 $\pm$ 0.02 0.02
w/o $\bm{w}^{sim}$ 0.88 $\pm$ 0.03 0.92 $\pm$ 0.02 0.87 $\pm$ 0.04 0.82 $\pm$ 0.04 0.17
w/o $\bm{w}^{per}$ 0.76 $\pm$ 0.06 0.88 $\pm$ 0.03 0.82 $\pm$ 0.05 0.87 $\pm$ 0.02 0.28
w/o both 0.63 $\pm$ 0.05 0.83 $\pm$ 0.04 0.63 $\pm$ 0.08 0.80 $\pm$ 0.05 1.24
Algo. 3DPickup Var. $\downarrow$ (×$\times 10^{-2}$)
Ball Cube Key Health kit
CenRA 0.95 $\pm$ 0.02 0.68 $\pm$ 0.09 0.80 $\pm$ 0.06 0.69 $\pm$ 0.07 1.57
w/o $\bm{w}^{sim}$ 0.82 $\pm$ 0.07 0.70 $\pm$ 0.09 0.70 $\pm$ 0.07 0.89 $\pm$ 0.04 0.89
w/o $\bm{w}^{per}$ 0.78 $\pm$ 0.07 0.40 $\pm$ 0.09 0.63 $\pm$ 0.08 0.44 $\pm$ 0.10 3.05
w/o both 0.81 $\pm$ 0.07 0.46 $\pm$ 0.06 0.48 $\pm$ 0.08 0.37 $\pm$ 0.08 3.80
Algo. MujocoCar Var. $\downarrow$ (×$\times 10^{3}$)
Target #1 Target #2 Target #3 Target #4
CenRA 588.2 $\pm$ 0.73 549.3 $\pm$ 0.64 447.7 $\pm$ 0.67 474.3 $\pm$ 0.66 4.24
w/o $\bm{w}^{sim}$ 319.9 $\pm$ 0.59 486.8 $\pm$ 0.71 332.5 $\pm$ 0.70 260.9 $\pm$ 0.54 9.29
w/o $\bm{w}^{per}$ 320.9 $\pm$ 0.89 355.3 $\pm$ 0.68 344.1 $\pm$ 0.87 308.2 $\pm$ 0.72 0.46
w/o both 57.5$\pm$0.35 257.0 $\pm$ 0.68 677.3 $\pm$ 0.54 77.6 $\pm$ 0.26 82.7
The results indicate that the two weights, which control the allocation of samples drawn from each policy agent’s experiences, mainly influence the overall learning performance. Specifically, the absence of sampling weight leads to unbalanced learning outcomes, which is observed by the increased variance in episodic returns across four tasks. While the full CenRA model does not always achieve the lowest variance, it consistently outperforms the other three ablation models regarding overall system performance.
Both weights play essential roles in the information synchronization, with the performance weight $\bm{w}^{per}$ having a more important impact. It allows CRA to pay more attention to the policy agents that are underperforming or progressing slowly. This suggests that considering the overall learning performance of the multi-task system is also an important objective that CenRA seeks to achieve.

### v3

@@ 4.1.3 Policy Agents with Knowledge Rewards
Each policy agent $\mathcal{A}^{pol}_{i}$ stores the experiences in its corresponding replay buffer $\mathcal{D}_{i}$. They receive two types of rewards: the environmental reward $r^{env}_{i}$ from their respective task $T_{i}$ and the knowledge reward $r^{knw}$ from CRA. The augmented reward is given by:
@@ 4.2 Information Synchronization of Policy Agents
In the CenRA, the information provided by different tasks may exhibit significant disparities, potentially leading to an imbalance in knowledge extraction and distribution. We introduce an information synchronization mechanism for CenRA to maintain a balanced manner from the perspective of the entire system. Specifically, we control the quantity of samples that CRA retrieves from each task’s replay buffer $\mathcal{D}_{i}$ by a sampling weight $\bm{w}$, by considering two aspects: the similarity among tasks and the real-time learning performance of the policy agents.
Similarity Weight $\bm{w}^{sim}$ is derived from the similarity among tasks, enabling the CRA to focus on relatively outlier tasks. To simplify computation, we use the hidden layers extracted from each policy agent’s neural network encoders to represent the tasks’ features. To reduce randomness, we average the hidden features of the most recent $K$ steps. We adopt a cross-attention mechanism to calculate the similarity weight [Vaswani et al., 2017]. Specifically, for task $T_{i}$, let $\bm{H}_{i}$ denote the averaged hidden feature vector, which serves as the key, and the centroid of all tasks $\bm{c}$ acts as the query. Then, the similarity $s_{i}$ of task $T_{i}$ to the centroid of the task cluster is calculated as:
@@
where $D$ is the dimension of the hidden feature to prevent gradient vanishing or exploding. A larger $s_{i}$ indicates a greater similarity between $T_{i}$ and the centroid. It is worth noting that, to avoid the centroid $\bm{c}$ approaching zero due to feature vectors $\bm{H}_{i}$ pointing in opposite directions, all latent representations $\bm{H}_{i}$ in our framework are extracted from ReLU activation layers. This ensures that every element of $\bm{H}_{i}$ is non-negative, effectively preventing feature cancellation and maintaining a well-defined and numerically stable centroid $\bm{c}$. Given our assumption is that the tasks farther from the centroid require more attention, the similarity weight is defined as $\bm{w}^{sim}=\text{Softmax}\big([1/s_{1},1/s_{2},\dots,1/s_{N}]\big)$.
@@
The final sampling weight $\bm{w}$ is formulated as $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$, where $\alpha$ is a hyperparameter to balance the two aspects. The CRA samples from each replay buffer $\mathcal{D}_{i}$ according to $\bm{w}$, ensuring a balanced and effective knowledge extraction and learning.
@@ 4.3 Overall Framework
The learned CRA acts as a robust knowledge pool, which is able to support new tasks by transferring knowledge through auxiliary reward signals. This is particularly beneficial in sparse-reward environments, as the knowledge rewards can guide the policy agents toward the correct direction and reduce exploration burden. Additionally, the CRA can be further optimized alongside new tasks in a continuous learning scheme that enhances adaptability and effectiveness in dynamic settings.
@@
3:   $(s_{t},a_{t},r^{env}_{t},s_{t+1},a_{t+1})\sim\text{Interact}(\mathcal{A}^{pol}_{i},T_{i})$ $\triangleright$ Interact and collect one transition
4:   $r^{knw}_{t}\sim\mathcal{A}^{rwd}(s_{t},a_{t})$ $\triangleright$ Sample an off-policy knowledge reward
@@
6:   $\mathcal{D}_{i}\leftarrow\mathcal{D}_{i}\cup\{(s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i})\}$ $\triangleright$ Store the transition in corresponding $\mathcal{D}_{i}$
7:   Update policy agent $\mathcal{A}^{pol}_{i}$ $\triangleright$ Update $\mathcal{A}^{pol}_{i}$ using backbone RL algorithm
@@
9:  $\bm{w}=\alpha\bm{w}^{sim}+(1-\alpha)\bm{w}^{per}$ $\triangleright$ Calculate sampling weight
10:  $\{s^{rwd}_{t},r^{knw}_{t},r^{env}_{t},s^{rwd}_{t+1}|T_{i}\}_{\mathcal{B}}\sim\mathcal{D}|\bm{w}$ $\triangleright$ Draw samples based on the sampling weight
@@ 5 Experiments
We conduct experiments in four MTRL domains as shown in Figure 2: the widely used Meta-World benchmark (including ML10 with 10 tasks and ML50 with 50 tasks) [Yu et al., 2020b], 2DMaze, 3DPickup [Chevalier-Boisvert et al., 2024], and MujocoCar [Ji et al., 2023]. All tasks, including those in Meta-World, are crafted to provide sparse environmental rewards, where the agent receives a reward of $1$ only upon successful completion of the final objective, and 0 otherwise. The detailed task configurations are provided in Appendix A.
@@
Figure 2: Environments with multiple tasks. (a) Meta-World: two sparse-reward versions are used: ML10-sparse and ML50-sparse, including diverse robotic manipulation tasks. (b) 2DMaze: 2D maze tasks where the agent must pick up a key and then pass through a door to exit. (c) 3DPickup: 3D maze tasks where the agent aims to navigate to and pick up different target objects at different locations. (d) MujocoCar: mujoco-based race car aims to navigate to different specified areas.
@@ 5.1 Comparative Evaluation in MTRL
We benchmark CenRA against several state-of-the-art baselines: (a) the backbone RL algorithms of the policy agents: DQN [Mnih et al., 2015] for discrete control tasks and SAC [Haarnoja et al., 2018b] for continuous control tasks; (b) the ReLara algorithm [Ma et al., 2024a], which can be regarded as a decentralized variant of CenRA, where each policy agent is paired with a separate reward agent, without cross-task information sharing; (c) the TD-MPC2 algorithm [Hansen et al., 2024]; (d) the Contrastive Modules with Temporal Attention (CMTA) algorithm [Lan et al., 2023]; (e) the Policy Optimization and Policy Correction (PiCor) algorithm [Bai et al., 2023]; (f) the Multi-Critic Actor Learning (MCAL) algorithm [Mysore et al., 2022]; (g) the Parameter-compositional MTRL (PaCo) algorithm [Sun et al., 2022]; (h) the Shared-Critic (SC) algorithm [Zhang et al., 2021]; and (i) the MTRL with Soft Modularization (SoftModule) [Yang et al., 2020]. They are implemented by either the CleanRL library [Huang et al., 2022] or official codebases. Each task is trained with 10 different random seeds, and the average results are reported.
@@ 5.2 Knowledge Transfer to New Tasks
In this section, we assess the CRA’s ability to transfer previously learned knowledge to unseen tasks. Specifically, we continue using the trained CRA model in Section 5.1, while initializing new policy agents to tackle new tasks from the same domain. These include 5 test tasks for ML10-sparse and ML50-sparse, and 1 test task for each of the remaining domains, none of which were encountered during the initial training. For the CenRA, we explore two scenarios: (1) the CRA continues to be optimized in collaboration with the new policy agent (CenRA w/ learning); and (2) only the policy agent is updated while the CRA remains fixed, relying only on its previously acquired knowledge (CenRA w/o learning). We compare the two settings against the backbone algorithms and ReLara. In ReLara, the reward agent is trained anew without pre-learned knowledge. The results are presented in Figure 4(a) and Table 2.
@@
(a) Comparison of the learning performance of CenRA with the baselines in new tasks in the 2DMaze, 3DPickup and MujocoCar domains.
(b) Actions yielding the highest knowledge rewards in a new 2DMaze task.
Figure 4: Experimental results for knowledge transfer to new tasks.
@@
To further demonstrate CenRA’s transferability, we select the 2DMaze environment to visualize the knowledge provided by CRA when facing an unseen task. As shown in Figure 4(b), we plot the directions of actions that yield the maximum knowledge reward at each position, categorized into two scenarios: before and after obtaining the key. While some guidance in peripheral regions may appear slightly misaligned, most states receive reasonable rewards that align with human understanding. This demonstrates the effectiveness of knowledge transfer and with such dense rewards, the agent’s adaptation to new tasks is able to be well supported. In addition, to better understand the knowledge learned by the CRA, we provide a case study in Appendix C, where we visualize the CRA-provided rewards. This further verifies that the CRA is capable of capturing domain-relevant, task-specific, and semantically meaningful signals across different tasks.
Algo. 2DMaze Var. $\downarrow$ ($\times 10^{-2}$)
Maze #1 Maze #2 Maze #3 Maze #4
CenRA ($\alpha=0.5$) 0.893 $\pm$ 0.033 0.908 $\pm$ 0.022 0.924 $\pm$ 0.020 0.932 $\pm$ 0.020 0.021
CenRA ($\alpha=0.25$) 0.889 $\pm$ 0.031 0.901 $\pm$ 0.025 0.915 $\pm$ 0.023 0.925 $\pm$ 0.024 0.065
CenRA ($\alpha=0.75$) 0.891 $\pm$ 0.030 0.905 $\pm$ 0.024 0.918 $\pm$ 0.021 0.928 $\pm$ 0.022 0.049
w/o $\bm{w}^{sim}$ ($\alpha=0$) 0.884 $\pm$ 0.033 0.922 $\pm$ 0.021 0.873 $\pm$ 0.041 0.820 $\pm$ 0.039 0.172
w/o $\bm{w}^{per}$ ($\alpha=1$) 0.758 $\pm$ 0.062 0.884 $\pm$ 0.030 0.824 $\pm$ 0.052 0.867 $\pm$ 0.020 0.284
w/o both 0.632 $\pm$ 0.053 0.833 $\pm$ 0.041 0.629 $\pm$ 0.08 0.802 $\pm$ 0.054 1.235
Algo. 3DPickup Var. $\downarrow$ ($\times 10^{-2}$)
Ball Cube Key Health kit
CenRA ($\alpha=0.5$) 0.951 $\pm$ 0.020 0.683 $\pm$ 0.090 0.795 $\pm$ 0.062 0.688 $\pm$ 0.067 1.570
CenRA ($\alpha=0.25$) 0.942 $\pm$ 0.023 0.671 $\pm$ 0.091 0.782 $\pm$ 0.066 0.715 $\pm$ 0.061 1.650
CenRA ($\alpha=0.75$) 0.938 $\pm$ 0.025 0.665 $\pm$ 0.095 0.775 $\pm$ 0.069 0.702 $\pm$ 0.065 1.723
w/o $\bm{w}^{sim}$ ($\alpha=0$) 0.822 $\pm$ 0.065 0.702 $\pm$ 0.093 0.704 $\pm$ 0.072 0.887 $\pm$ 0.038 0.892
w/o $\bm{w}^{per}$ ($\alpha=1$) 0.779 $\pm$ 0.072 0.404 $\pm$ 0.093 0.631 $\pm$ 0.080 0.438 $\pm$ 0.102 3.051
w/o both 0.811 $\pm$ 0.073 0.457 $\pm$ 0.058 0.483 $\pm$ 0.079 0.370 $\pm$ 0.079 3.796
Algo. MujocoCar Var. $\downarrow$ ($\times 10^{3}$)
Target #1 Target #2 Target #3 Target #4
CenRA ($\alpha=0.5$) 588.221 $\pm$ 0.732 549.337 $\pm$ 0.640 447.743 $\pm$ 0.672 474.320 $\pm$ 0.657 4.244
CenRA ($\alpha=0.25$) 575.153 $\pm$ 0.740 538.912 $\pm$ 0.655 439.850 $\pm$ 0.680 462.116 $\pm$ 0.665 4.871
CenRA ($\alpha=0.75$) 580.431 $\pm$ 0.735 542.765 $\pm$ 0.648 441.033 $\pm$ 0.675 468.529 $\pm$ 0.660 4.533
w/o $\bm{w}^{sim}$ ($\alpha=0$) 319.926 $\pm$ 0.590 486.767 $\pm$ 0.712 332.506 $\pm$ 0.695 260.921 $\pm$ 0.544 9.288
w/o $\bm{w}^{per}$ ($\alpha=1$) 320.887 $\pm$ 0.891 355.325 $\pm$ 0.677 344.145 $\pm$ 0.872 308.215 $\pm$ 0.723 0.457
w/o both 57.532 $\pm$ 0.352 257.010 $\pm$ 0.677 677.285 $\pm$ 0.540 77.635 $\pm$ 0.255 82.720
@@ 5.3 Effect of Sampling Weight
We conduct experiments to understand the effects of the information synchronization mechanism in the CenRA. Specifically, we compare the full CenRA model against five variants: (a) and (b) CenRA with different values of the balance factor $\alpha$, i.e., $\alpha=0.25$ and $\alpha=0.75$, to examine the impact of different weight combinations; (c) CenRA without the similarity weight $\bm{w}^{sim}$ (i.e., $\alpha=0$); (d) CenRA without the performance weight $\bm{w}^{per}$ (i.e., $\alpha=1$); and (e) CenRA without the entire sampling weight. To better illustrate the differences among tasks and highlight the role of sampling weights in task coordination and synchronization, we select the four-task domains, i.e., 2DMaze, 3DPickup, and MujocoCar. The results are shown in Table 3.
The results indicate that the two weights, which control the allocation of samples drawn from each policy agent’s experiences, mainly influence the overall learning performance. Specifically, the absence of sampling weight leads to unbalanced learning outcomes, which is observed by the increased variance in episodic returns across four tasks. In contrast, when both weights are incorporated, the learning process becomes notably more stable, indicating that the joint consideration of task similarity and learning progress is essential for coordinated optimization. While the full CenRA model does not always achieve the lowest variance, it consistently outperforms the other three ablation models regarding overall system performance.
Both weights play essential roles in information synchronization, with the performance weight $\bm{w}^{\mathrm{per}}$ having a more significant impact. It allows the CRA to focus more on policy agents that are underperforming or progressing slowly, ensuring balanced system-wide learning. Moreover, different choices of the balancing factor $\alpha$ emphasize distinct aspects of synchronization: a larger $\alpha$ highlights task similarity and promotes uniformity across related tasks, whereas a smaller $\alpha$ prioritizes compensating lagging tasks by amplifying the effect of $\bm{w}^{\mathrm{per}}$. This flexible weighting further enhances stability and adaptability, demonstrating that considering the overall learning performance of the multi-task system is a central objective that CenRA seeks to achieve.